Find the slope and the y-intercept of the line.
2y+4x+5=20+4x-2y

Answer by haileytucki(390)

2y+4x+5=20+4x-2y
Since -2y contains the variable to solve for, move it to the left-hand side of the equation by adding 2y to both sides.
2y+4x+5+2y=4x+20
Since 2y and 2y are like terms, add 2y to 2y to get 4y.
4y+4x+5=4x+20
Move all terms not containing y to the right-hand side of the equation.
4y=-4x-5+4x+20
Simplify the right-hand side of the equation.
4y=15
Divide each term in the equation by 4.
(4y)/(4)=(15)/(4)
Simplify the left-hand side of the equation by canceling the common factors.
y=(15)/(4)
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula, m=0 and b=(15)/(4).
m=0, b=(15)/(4)